There are four steps to algorithm methodology.

1. Design: Identify the problem and thoroughly understand it.

2. Analyze how efficient the code is in solving the problem

3. Implement: Writing and coding the algorithm.

4. Experiment with different variables in the algorithm.

- Use Python3 to create your own SHORT algorithm to complete a computing task.

Discuss your algorithm design in relation to these four steps and describe how you went through each step of the methodology to create your algorithm.
Answer #1

Design: The problem is to sort a unsorted array, this can be done in either compare each element of the array to the other elements of the array and placing them in the right place but it will take lots of time as two loops will be required for it. So efficient algo can be:

  1. Take a array
  2. for i=0 to array length
  3. key <= arr[i]
  4. j <= i-1
  5. while (j >= 0 && arr[j] > key)
  6. arr[j + 1] = arr[j];
  7. j <= j - 1;
  8. end
  9. arr[j+1]=key
  10. end

Analyse: O(n^2)

Implement:

def sort(arr):
   for i in range(1, len(arr)):
       key = arr[i]
       j = i-1
       while j >= 0 and key < arr[j] :
               arr[j + 1] = arr[j]
               j = j-1
       arr[j + 1] = key